# A Mathematical Definition of Path Integrals on Symplectic Manifolds

### v1

@@ Abstract
We give a mathematical definition of some path integrals, emphasizing those relevant to the quantization of symplectic manifolds (and more generally, Poisson manifolds) — in particular, the coherent state path integral. We show that Kähler manifolds provide many computable examples.
@@ 1 The Simplest Case
- 4 More Examples From Kähler Manifolds
- 5 Relation to Geometric Quantization
@@ 0 Introduction
for In eq. 0.0.1, is the connection on the prequantum line bundle and the integrand should be interpreted as parallel transport from the vector space over to the vector space over 222Implicit in this notation is a choice of orthonormal bases of the vector spaces over We have written it this way to be consistent with the physics literature, eg. equation (2.7) of [9]. See section 2 for the definition. Computing this path integral immediately produces a quantization, ie. a good Hilbert space and quantization map defined on all classical observables, as explained in section 1, section 2. We suggest that computing it is equivalent to finding a propagator satisfying definition 2.0.1. In eq. 0.0.2, the integral is over maps from the disk into where are points on the boundary of the disk. Equation 0.0.2 naturally generalizes to Poisson manifolds. Kähler manifolds produce many computable examples, and as such this is related to Toeplitz quantization [3], as discussed in [5].
@@
The idea is that we can compute eq. 0.0.1 by finding a function on with the right derivative. We will not prove any very general existence theorems, but we do obtain well-defined existence and uniqueness questions. This paper is largely a continuation of [18], where we formalized the procedure of putting a path integral on a lattice, in a wide class of cases (eg. for path integrals over spaces of maps between manifolds). That paper describes a path integral construction analogous to the construction of the Riemann integral via Riemann sums, eg. Feynman’s construction of the path integral. This paper, on the other hand, focuses on the path integral analogue of the fundamental theorem of calculus.
@@ 1.1 Definition of the Path Integral
be a smooth function, considered as an integral kernel with respect to Suppose that satisfies the following four conditions (compare with eq. 0.0.4, 0.0.5): for all
The first condition says that the convolution and this is one of the defining properties of a (time-independent) propagator;444See remark 1.1.2 for a comment on the time-dependent case. is the exterior derivative in the second factor; the additional third and fourth conditions are normalization and reversal of orientation properties.555These condition is automatically satisfied by where satisfies eq. 0.0.4 and with the multiplicative structure replaced by the additive structure. Together with eq. 1.1.2, they imply that
@@
Now, if we ignore condition eq. 1.1.2, then the limit of the right side of eq. 1.1.9 is essentially Feynman’s method of defining eq. 1.1.11, which is made precise by the lattice construction given in [18]. Formally then, the right side of eq. 1.1.9 is a Riemann sum with the algebraic property of eq. 0.0.4 replaced by eq. 1.1.2 (of course, we have to understand Riemann sums more generally to make this statement, see section 3).
By the complex version of Komolgorov’s extension theorem (A.2 of [21]), such an defines a complex measure on the space of paths though a priori this measure is concentrated on all paths (including discontinuous ones).777One should be able to modify the measurable space so that this measure is concentrated on a smaller set of discontinuous paths. A relevant discussion is in Berezin [1]. We can denote this measure by
@@ 1.2 The Hilbert space and Path Integral – Operator Correspondence
On eq. 1.2.1 agrees with the -inner product. This eigenspace is analogous to a space of complex polarized sections of a prequantum line bundle, and we can identify it with where if
Such an is sometimes called a reproducing kernel.101010Though this Hilbert space is not quite a reproducing kernel Hilbert space since it doesn’t contain all functions on has the special property that pointwise evaluation is a bounded linear functional, since by the definition of for
@@
To be clear, the correspondence between the path integral and the operator formulations is an instance of the correspondence determined by the Riesz representation theorem between sesquilinear forms and operators.111111However, there is a sense in which the path integral formulation logically precedes the operator formalism. This is also one reason why the coherent space path integral (or phase space path integral, [14]) is nice, as this correspondence is obfuscated when a polarization is used.
@@ 1.3 Example: Conventional Quantum Mechanics
The state is identified with the corresponding quantum mechanical coherent state, ie. the eigenvector of the lowering operator121212These are usually defined in the context of the harmonic oscillator. for which the expectation values of the momentum and position operators are respectively. The physical Hilbert space is the Segal-Bargmann representation (sometimes called the holomorphic representation), [13].
Many other examples come from Kähler manifolds, as we will discuss in section 4.
@@ 1.4.1 Special Observables
On the left is a quantum expectation value and on the right is a classical expectation value.131313These describe states with classical statistical uncertainty. In [6], the dual quantization scheme is used, ie. a function quantizes to an operator such that However, such a quantization map doesn’t need to exist.
@@ 1.4.2 Deformation Quantization
In this case we obtain a formal deformation quantization, see equation (2.10) of [9]. We claim that this is what occurs in Toeplitz quantization [3]. We will explore this in forthcoming work, also see section 2.1 for more on this relation.
@@ 2 The General Case: The Propagator of Symplectic Manifolds
More generally, we can replace with square-integrable sections of a complex line bundle with Hermitian connection, and essentially everything still works. Of course, the most natural setting for this is a prequantizable symplectic manifold which comes with a natural measure. Even better are Kähler manifolds, since the path integral can be computed more canonically in this case due to it coming with a Hermitian form, as discussed in section 4.1.
In this setting, consider a prequantum line bundle with Hermitian connection
where we assume In this case, the path integral we want to compute is
where denotes parallel transport between the vector spaces over and 141414We have supressed the dependence, but it divides That is, eq. 2.0.2 must define a linear map between any two vector spaces of the line bundle. The kernel we need to consider then is a section of
@@
The conditions of definition 1.1.1 are replaced by: for all 151515In this notation, we are assuming acts on the left, but we may not be consistent. Furthermore, for the first condition it is acceptable to scale by a finite constant.
To explain eq. 2.0.5, we note that each defines a section of given by
where the subscript is used to denote the corresponding vector space. Equation 2.0.5 means that the pushforward of this map agrees with as maps for all
@@
We discuss uniqueness of such a propagator in section 4.1.
@@
Therefore, associated to there is a canonical projection operator, which we denote by It follows that
@@
Interestingly, if we can find a subspace of sections of the prequantum line bundle for which pointwise evaluation is continuous, then we can try to flip this construction to determine a propagator via eq. 2.0.14. This leads us to our next example.
Example 2.0.3.
Let be a prequantizable Kähler manifold, with prequantum line bundle
Then pointwise evaluation of complex-polarized sections is a continuous map, and we can define by
where is defined via the Riesz representation theorem, as above. is a propagator integrating and the eigenspace of the corresponding operator consists of complex-polarized sections. See [5] for more.
Note that, the previous example somewhat contradicts the discussion in section 2.3 of [9]. This is due to the fact that this path integral isn’t invariant under all symplectomorphisms, as is assumed.
Lemma 2.0.4.
181818See appendix .1 for details about the van Est map,If is a propagator integrating then
where is interpreted as a function and is now the symplectic form.
In the previous lemma, should be interpreted as
where the domain of integration is all maps with three marked points on mapping to (compare with equations (2.9), (2.11) of [9]).
@@ 2.1 Relation to Kontsevich’s Work, the Poisson Sigma Model and Weinstein’s Program
The propagator of the previous subsection defines a section of the prequantum line bundle over the symplectic groupoid, and it is an idempotent with respect to the twisted convolution algebra ([11]). Equation 2.0.5 implies that where is the Poisson tensor and is the groupoid pullback (this makes sense because of this prequantum line bundle is canonically trivial). Therefore, on a Poisson manifold this construction gives a -algebra and quantization map, as in Weinstein’s program [26], and can be used to compute
Here,
and the path integral is over all Lie algebroid morphisms where are marked points on This is formally equivalent to the Poisson sigma model approach to Kontsevich’s star product [15], see [2]. In the symplectic case, it is the same as eq. 0.0.2. We will explore this more in forthcoming work.
@@ 3.3 Differentiating Path Measures to Differential Forms
4 More Examples From Kähler Manifolds
@@
Let be a prequantizable Kähler manifold, with prequantum line bundle
Then as in example 2.0.3, pointwise evaluation of complex-polarized sections is a continuous map. Given a normalized vector over we get a continuous linear functional on complex-polarizared sections given by
where is defined by The Riesz representation theorem determines a complex-polarized section corresponding to this linear functional. We can define by
as in lemma 2.0.4. This is well-defined, ie. it is independent of the choices of and
If is simply connected, then should be interpreted as
where the path integral is over all maps such that where are three marked points on the boundary of the disk. See eq. 0.0.2.
@@ 4.1 Surfaces of Constant Curvature
Now we will recall some examples of eq. 0.0.1 computed by Daubechies, Klauder, on Kähler manifolds in [6], [12] (see [5] for the general Kähler case). We will show that they satisfy definition 2.0.1. To arrive at these expressions, the aforementioned authors (essentially) introduce Brownian motion into eq. 0.0.1 via a massive term and send
@@
The following locally describe some 2-dimensional Kähler manifolds and their propagators eq. 0.0.1, for which the connection on the prequantum line bundle is and is the Riemannian metric (which has constant curvature).
@@
Hyperbolic plane: 222222The examples computed by Klauder depend on parameters which we’ve set equal to
@@
(where the exterior derivative is taken in Therefore, integrates in the sense of definition 2.0.1. This illustrates the non-uniqueness of the value of eq. 0.0.1.
Let us emphasize the following: there is a unique solution to eq. 0.0.4, 0.0.5, while the corresponding solution to eq. 1.1.2-1.1.5 (or eq. 2.0.4-2.0.7) is not unique, and this is for the same reason that the Riemann integral doesn’t have a unique value under path integrals, as discussed in section 3.2 — in order to produce a unique value, one needs to specify a second order term, as in eq. 3.2.5.
More interesting then, is the second order Taylor expansion in the variables of at This does distinguish between the different propagators. To make sense of this, we use the corresponding Riemannian metric to split the short exact sequence associated to the jet bundles
@@
We might conjecture that this is true for general Kähler manifolds. Geometrically, this is equal to
@@
with the understanding that eq. 4.1.12 doesn’t hold on non-smooth paths. It would be interesting to know if the second order Taylor expansions (or 2-jets) uniquely determine the propagator — these 2-jets are given by a higher order van Est map ([18]). Similarly, it would be interesting to know if the lattice approach in [18] converges to the propagator. For example, we could cover with small enough neighborhoods such that there is a unique geodesic connecting any two points in such a neighborhood, and approximate the integrand of eq. 4.1.13 using
where is the distance between the corresponding points, and is the unique geodesic in the neighborhood connecting We can then compute the path integral as in Feynman’s approach, only integrating in such a way that are in the same neighborhood.
5 Relation to Geometric Quantization
In section 1.3 we discussed the relation of these propagators to geometric quantization, using complex polarizations (also see section 2.1). However, one can ask about real polarizations. In [20] (also see [19]) we discussed a function

### v2

@@ Abstract
We give a mathematical definition of some path integrals, emphasizing those relevant to the quantization of symplectic manifolds (and more generally, Poisson manifolds) — in particular, the coherent state path integral. We show that Kähler manifolds provide many computable examples and we emphasize those whose Bergman kernel is constant along the diagonal.
@@ 1 The Simplest Case
- 4 Examples
- 5 Relation to Other Quantization Schemes
@@ 0 Introduction
for In eq. 0.0.1, is the connection on the prequantum line bundle and the integrand should be interpreted as parallel transport over from the vector space over to the vector space over 222Implicit in this notation is a choice of orthonormal bases of the vector spaces over We have written it this way to be consistent with the physics literature, eg. equation (2.7) of [11]. See section 2 for the definition. Computing this path integral immediately produces a quantization, ie. a good Hilbert space and quantization map defined on all classical observables, as explained in section 1, section 2. We will call it a propagator.
We suggest that computing eq. 0.0.1 is equivalent to finding a propagator satisfying definition 2.0.1. In eq. 0.0.2, the integral is over maps from the disk into where are points on the boundary of the disk. It is related to eq. 0.0.1, since
Equation 0.0.2 naturally generalizes to Poisson manifolds. Kähler manifolds produce many computable examples, with particularly nice ones arising from those whose Bergman kernel is constant along the diagonal. As such this is closely related to Toeplitz quantization [3], as discussed in [6], [28].
@@
The idea is that we can compute eq. 0.0.1 by finding a function on with the right derivative. We will not prove any very general existence theorems, but we do obtain well-defined existence and uniqueness questions. This paper is largely a continuation of [22], where we formalized the procedure of putting path integrals over morphisms of Lie algebroids on a lattice (eg. path integrals over maps between manifolds, section 4.2.2). That paper describes a path integral construction analogous to the construction of the Riemann integral via Riemann sums, eg. Feynman’s construction of the path integral. This paper, on the other hand, focuses on the path integral analogue of the fundamental theorem of calculus — the construction of the propagator fits into the framework of the aforementioned paper.
In particular, we have the following lemma:
Lemma.
Indeed, the examples computed by Daubechies, Klauder all have the property that the Bergman kernel is constant along the diagonal. The significance in geometric quantization of it being constant has been known at least since [4] (where they also give a sufficient condition for it to be constant), and such prequantizations were studied by Donaldson in [8], where they are called balanced. In general, the normalized Bergman kernel integrates the prequantum connection of the Bergman Kähler form, which is an -perturbation of the original Kähler form. It is a rescaling of the pullback of the Fubini-Study form by the Kodaira embedding, [26], [33].
@@ 1.1 Definition of the Path Integral
be a smooth function, considered as an integral kernel with respect to Suppose that satisfies the following three conditions (compare with eq. 0.0.5, 0.0.6): for all
The first condition says that the convolution and this is one of the defining properties of a (time-independent) propagator;444See remark 1.1.2 for a comment on the time-dependent case. is the exterior derivative in the second factor; the additional third and fourth conditions are normalization and reversal of orientation properties.555These condition is automatically satisfied by where satisfies eq. 0.0.5 and with the multiplicative structure replaced by the additive structure.
@@
Now, if we ignore condition eq. 1.1.2, then the limit of the right side of eq. 1.1.8 is essentially Feynman’s method of defining eq. 1.1.10, which is made precise by the lattice framework developped in [22], see section 4.2.2. Formally then, the right side of eq. 1.1.8 is a Riemann sum with the algebraic property of eq. 0.0.5 replaced by eq. 1.1.2 (of course, we have to understand Riemann sums more generally to make this statement, see section 3).
By the complex version of Komolgorov’s extension theorem (A.2 of [25]), such an defines a complex measure on the space of paths though a priori this measure is concentrated on all paths (including discontinuous ones).777One should be able to modify the measurable space so that this measure is concentrated on a smaller set of discontinuous paths. A relevant discussion is in Berezin [1]. In the case of a non-trivial line bundle, as in the next section, the measure is operator-valued. We can denote this measure by
@@ 1.2 The Hilbert space and Path Integral – Operator Correspondence
On eq. 1.2.1 agrees with the -inner product. This eigenspace is analogous to a space of Kähler-polarized sections of a prequantum line bundle, and we can identify it with where if
Such are sometimes called reproducing kernels and have been widely studied, eg. [15], [29]. has the special property that pointwise evaluation is a bounded linear functional, since by the definition of for
@@
To be clear, the correspondence between the path integral and the operator formulations is an instance of the correspondence determined by the Riesz representation theorem between sesquilinear forms and operators.101010However, there is a sense in which the path integral formulation logically precedes the operator formalism. This is also one reason why the coherent space path integral (or phase space path integral, [18]) is nice, as this correspondence is obfuscated when a polarization is used.
@@ 1.3 Example: Conventional Quantum Mechanics
The state is identified with the corresponding quantum mechanical coherent state, ie. the eigenvector of the lowering operator111111These are usually defined in the context of the harmonic oscillator. for which the expectation values of the momentum and position operators are respectively. The physical Hilbert space is the Segal-Bargmann representation (sometimes called the holomorphic representation), [17].
Many other examples come from Kähler manifolds, as we will discuss in section 4.2.
@@ 1.4.1 Special Observables
On the left is a quantum expectation value and on the right is a classical expectation value.121212These describe states with classical statistical uncertainty. In [7], the dual quantization scheme is used, ie. a function quantizes to an operator such that However, such a quantization map doesn’t need to exist.
@@ 1.4.2 Deformation Quantization
In this case we obtain a formal deformation quantization, see equation (2.10) of [11]. This is what occurs in Toeplitz quantization [3]. See section 5.1, [28] for more on this relation. To make sense of what it measn to be an isomorphism perturbatively, we give the following definition:
Definition 1.4.1.
(see [14]) Let be a Poisson manifold and let be a set containing as an accumulation point. For each let be a unital -algebra such that and let
be ∗- linear131313This compatibility with ∗ can be relaxed. such that its image generates Furthermore, assume that is the inclusion map. We say that is a (non-perturbative) deformation quantization of if there is a star product on such that, for all
where is the component of the star product up to order
In this definition, we don’t have to assume that is a star product a priori, just that it is an asymptotic expansion in
@@ 2 The General Case: The Propagator of Symplectic Manifolds
More generally, we can replace with square-integrable sections of a complex line bundle with Hermitian connection, where of the previous section corresponds to the connection, and everything essentially works the same. Of course, the most natural setting for this is a prequantizable symplectic manifold which comes with a natural measure. Even better are Kähler manifolds, since there are canonically defined propagators, as discussed in section 4.
In this setting, consider a line bundle with Hermitian connection
In this case, the path integral we want to compute is
where denotes parallel transport between the vector spaces over and over the curve That is, eq. 2.0.2 must define a linear map between any two vector spaces of the line bundle. The kernel we need to consider then is a section of
@@
The conditions of definition 1.1.1 are replaced by: for all
The subscript in eq. 2.0.5 means that the covariant derivative is being taken in the second component; each defines a section of given by
where the subscript is used to denote the corresponding vector space. Equation 2.0.5 says that the covariant derivative of this section at is zero.
@@
If satisfies the previous conditions and the corresponding sesquilinear form eq. 2.0.19 is bounded, we say that it integrates (as a propagator).151515We may relax eq. 2.0.4, see section 5.
Any section of eq. 2.0.3 which equals on the diagonal determines a connection on (by differentiating in the second component at the diagonal),161616We are using the identification of splittings with connections. and eq. 2.0.5 says that this connection is equal to Furthermore, for any smooth path and triangulation of it follows that
converges to parallel transport over as We discuss uniqueness of such a propagator in section 4.2.
Remark 2.0.2.
In the symplectic case eq. 0.0.1, we can interpret the exponent as specifying the connection only to leading order in in which case it is natural to relax eq. 2.0.5 to
See remark 4.1.3.
The next lemma tells us that that if we can find such an without assuming eq. 2.0.6, then we can always normalize so that all conditions are satisfied:
Lemma 2.0.3.
Suppose satisfies the following conditions: for all
Then is a propagator integrating where is given by the following: let then
@@
Therefore, associated to there is a canonical projection operator, which we denote by 181818As pointed out to me by Francis Bischoff, the map is a version of the Kodaira embedding. In addition, it follows that
where is any vector over with unit norm.
@@
Interestingly, if we can find a subspace of sections of the prequantum line bundle for which pointwise evaluation is continuous, then we can try to flip this construction to determine a propagator via eq. 2.0.24. We will do this in section 4.1.
@@ 3.3 Differentiating Path Measures to Differential Forms
4 Examples
4.1 Toeplitz Quantization and Bergman Kernels
Let be a prequantizable Kähler manifold, with prequantum line bundle
Then pointwise evaluation of Kähler-polarized sections is a continuous map, and as previously discussed each unit-norm vector over determines a linear functional given by where
We denote by the associated section determined by the Riesz representation theorem. Note that, is independent of and we denote it by
We can define by
where have unit norm (this is independent of ).
The corresponding operator (section 1.2) is the orthogonal projection of onto the subspace of holomorphic sections. The quantization map determined by the path integral is equal to the Toeplitz quantization map [3]. Its corresponding integral kernel is given by
and this is polarized with respect to the Kähler polarization on 222222To be more clear, it is polarized with respect to a possibly different connection — it is this integral kernel multiplied by that is polarized with respect to (consistent with [13]).
is closely related to the Bergman kernel ([34]), which is important in the quantization of Kähler manifolds ([28]), and is called the Bergman measure (or coherent measure in [15]). The Bergman kernel can be described as follows: let be an orthonormal basis for the space of holomorphic sections. Then is a section of and is given by
We then define a propagator232323Note that, is canonically trivial over the diagonal, so we can identify with a positive-valued function.
Note that,
Lemma 4.1.1.
integrates with respect to the Bergman measure
if and only if is constant.
Proof.
We need to show that if and only if is constant on the diagonal. For the only if direction, differentiating along vectors in the eigenspace and using the product rule implies that is a holomorphic function. However, it is also real-valued, so it must be constant.
The other direction follows from product rule: since and is constant, for vectors we have that
Since is antiholomorphic in it follows that for in the eigenspace (ie. of the form Since it is also true for in the eigenspace (since is holomorphic in ) the result follows. ∎
@@ 4.1 Toeplitz Quantization and Bergman Kernels
is constant for where is the Fubini-Study symplectic form ([34]).
We get a function given by
This defines a function on composable arrows of the pair groupoid, and if is constant, then applying to eq. 5.1.2 gives This is closely related to deformation quantization. Equation 5.1.2 should be interpreted as
where the domain of integration is all maps with three marked points on mapping to (compare with equations (2.9), (2.11) of [11]). If is simply connected then
where is any map which restricts to on This corresponds to eq. 0.0.2 (after gauge reduction, see [2]).
For a relation of this quantization scheme to Berezin’s star product and the Poisson sigma model on Kähler manifolds, see [28].
Note that, these examples somewhat contradicts the discussion in section 2.3 of [11]. This is due to the fact that the propagator isn’t invariant under all symplectomorphisms, as is assumed.
Remark 4.1.3.
The map is always asymptotically constant as ([28]), and integrates a perturbed connection — the curvature of this perturbed connection is the Bergman Kähler form, which is an -perturbation of the original symplectic form, [26], [33]. Therefore, it’s still true that
which is what is needed for a quantization. See remark 2.0.2. When is constant we have an exact equality.242424The property of the Bergman kernel being constant was used in the context of Kähler quantization in [4]. For it to be constant, it is enough that the manifold and line bundle are homogeneous.
The function is a well-studied map, and it being constant has other interesting consequences, eg. it implies stability in GIT (see page 2 of [27]). Metrics such that this map is constant are called balanced, [8]. For a generalization of the Bergman kernel to symplectic manifolds, see section 4 of [29] and [15].
@@ 4.2 Surfaces of Constant Curvature
Here we will recall some explicit examples of the coherent state path integral eq. 0.0.1 computed by Daubechies, Klauder, on Kähler manifolds in [7], [16]. We will show that they satisfy definition 2.0.1. To arrive at these expressions, the aforementioned authors (essentially) introduce Brownian motion into eq. 0.0.1 via a massive term and send 252525For more explicit examples of Bergman kernels, see [15]. In all of the following, the kernels are constant along the diagonal. In this paper, we have argued that this is true in general: whenever the Bergman kernel is constant along the diagonal, it computes the coherent state path integral.
@@
The following locally describe some 2-dimensional Kähler manifolds and their propagators eq. 0.0.1 in Darboux coordinates, for which the connection on the prequantum line bundle is and is the Riemannian metric (which has constant curvature).
@@
Hyperbolic plane: 262626The examples computed by Klauder depend on parameters which we’ve set equal to
@@
(where the exterior derivative is taken in Therefore, integrates in the sense of definition 2.0.1. This illustrates the non-uniqueness of the value of eq. 0.0.1, since the complement of an arc connecting the north and south pole of the sphere (which has measure zero) is symplectomorphic to with its standard symplectic form.
Let us emphasize the following: there is a unique solution to eq. 0.0.5, 0.0.6, while the corresponding solution to eq. 1.1.2-1.1.5 (or eq. 2.0.4-2.0.6) is not unique, and this is for the same reason that the Riemann integral doesn’t have a unique value under path integrals, as discussed in section 3.2 — in order to produce a unique value, one needs to specify a second order term, as in eq. 3.2.5. This leads us into the following:
4.2.1 The Second Order Taylor Expansion of the Propagator (Bergman Kernel)
Since the first order Taylor expansion doesn’t uniquely determine the propagator, more interesting then is the second order Taylor expansion in the variables of at This does distinguish between the different propagators. To make sense of this, we use the corresponding Riemannian metric to split the short exact sequence associated to the jet bundles
@@ 4.2.1 The Second Order Taylor Expansion of the Propagator (Bergman Kernel)
Geometrically, this is equal to
@@
with the understanding that eq. 4.2.12 doesn’t hold on non-smooth paths. It would be interesting to know if the second order Taylor expansions (or 2-jets) uniquely determine the propagator — these 2-jets are given by a higher order van Est map ([22]).
4.2.2 The Lattice Approach to the Path Integral
Similarly, it would be interesting to know which path integral approximations given by the lattice approach of [22] converge to the propagator. This involves choosing a section of over a neighborhood of the diagonal in such that: restricted to the diagonal is equal to and its derivative in the second component, evaluated on the diagonal, agrees with the connection — this means that the connection determined by agrees with 272727We can always use the the Bergman kernel itself for , in which case the following lattice construction clearly works.
For example, we could cover with small enough open sets such that there is a unique geodesic connecting any two points in any of these open sets — this determines a neighborhood of the diagonal in with if for some We get a section of over with the correct derivative, given by
where is the distance between the corresponding points and denotes parallel transport over the geodesic connecting to For a map and for a triangulation of we get an approximation to the integrand of eq. 4.2.13 given by
We can then compute the path integral eq. 4.2.13 as in Feynman’s approach, where the measure is approximated using and the normalization is determined by the fact that the restriction of the path integral to the diagonal is equal to the That is, we can try to compute
where and is determined by the fact that for each the approximation is equal to if
Note that indeed, the derivative (or pushforward) of in the second component, evaluated at the diagonal, agrees with Classicaly, this is enough information to get a well-defined integral (see section 3) — the point is that, this is all of the information that is determined by the formal path integral. We can use the metric to split the short exact sequence associated to the jet bundle, eq. 4.2.8, and using this splitting also has the correct second derivative per eq. 4.2.13, ie.
5 Relation to Other Quantization Schemes
In the context of symplectic manifolds, the propagator is a section of the prequantum line bundle over the symplectic groupoid, and it is an idempotent with respect to the twisted convolution algebra ([13]). Therefore, definition 2.0.1 naturally generalizes to Poisson manifolds. Equation 2.0.5 implies that
where is the Poisson bivector and
is a function defined on composable arrows. To understand this function, consider three marked points on Equation 5.1.2 should be interpreted as
where the integral is over morphisms whose restrictions to
are algebroid-homotopic to respectively, and where is any morphism whose restriction to is 303030The symplectic groupoid being prequantizable implies that for morphisms hence this is well-defined. Therefore, on a Poisson manifold this construction gives a -algebra and quantization map, as in Weinstein’s program [36], and can be used to compute
Here, the integral is over all contractible morphisms and is any morphism whose restriction to is This is formally equivalent to the Poisson sigma model approach to Kontsevich’s star product [19], see [2]. In the symplectic case, it is the same as eq. 0.0.2 (after gauge reduction). See section 1.4.2.
5.2 Relation to Kostant-Souriau’s Geometric Quantization
In section 1.3 we discussed the relation of these propagators to geometric quantization, using Kähler polarizations (also see section 5.1). However, one can ask about real polarizations.313131Of course, there are examples where the propagator can be indirectly obtained using real polarizations, as discussed in section 1.3. In this section we discuss viewing real-polarized sections as the physical states associated to a propagator. In [24] (also see [23]) we discussed a function